# LLM Internal States Reveal Hallucination Risk Faced With a Query

### v1

@@ Appendix D AI Assistants Using
Question Answering 207 task837 viquiquad answer generation, task701 mmmlu answer generation high school computer science, task1399 obqa answer generation, task075 squad1.1 answer generation, task724 mmmlu answer generation moral scenarios, task666 mmmlu answer generation astronomy, task742 lhoestq answer generation frequency, task1438 doqa cooking answer generation, task863 asdiv multiop question answering, task864 asdiv singleop question answering, task058 multirc question answering, task669 ambigqa answer generation, task704 mmmlu answer generation high school government and politics, task728 mmmlu answer generation professional accounting, task740 lhoestq answer generation quantity, task1293 kilt tasks hotpotqa question answering, task849 pubmedqa answer generation, task1424 mathqa probability, task1625 disfl qa asnwer generation, task858 inquisitive span detection, task723 mmmlu answer generation moral disputes, task083 babi t1 single supporting fact answer generation, task118 semeval 2019 task10 open vocabulary mathematical answer generation, task582 naturalquestion answer generation, task237 iirc answer from subtext answer generation, task714 mmmlu answer generation human sexuality, task444 com qa question paraphrases answer generation, task720 mmmlu answer generation marketing, task332 tellmewhy answer generation, task119 semeval 2019 task10 geometric mathematical answer generation, task310 race classification, task1132 xcsr ur commonsense mc classification, task702 mmmlu answer generation high school european history, task710 mmmlu answer generation high school statistics, task870 msmarco answer generation, task047 miscellaneous answering science questions, task711 mmmlu answer generation high school us history, task1286 openbookqa question answering, task598 cuad answer generation, task685 mmmlu answer generation clinical knowledge, task084 babi t1 single supporting fact identify relevant fact, task1420 mathqa general, task1520 qa srl answer generation, task868 mawps singleop question answering, task768 qed text span selection, task061 ropes answer generation, task041 qasc answer generation, task144 subjqa question answering, task1570 cmrc2018 answer generation, task1610 xquad es answer generation, task164 mcscript question answering text, task703 mmmlu answer generation high school geography, task705 mmmlu answer generation high school macroeconomics, task1131 xcsr es commonsense mc classification, task1130 xcsr vi commonsense mc classification, task750 aqua multiple choice answering, task473 parsinlu mc classification, task385 socialiqa incorrect answer generation, task691 mmmlu answer generation college physics, task719 mmmlu answer generation management, task1327 qa zre answer generation from question, task715 mmmlu answer generation international law, task737 mmmlu answer generation world religions, task010 mctaco answer generation event ordering, task741 lhoestq answer generation place, task028 drop answer generation, task730 mmmlu answer generation professional medicine, task491 mwsc answer generation, task716 mmmlu answer generation jurisprudence, task732 mmmlu answer generation public relations, task735 mmmlu answer generation us foreign policy, task898 freebase qa answer generation, task887 quail answer generation, task024 cosmosqa answer generation, task1140 xcsr pl commonsense mc classification, task225 english language answer generation, task1608 xquad en answer generation, task170 hotpotqa answer generation, task667 mmmlu answer generation business ethics, task699 mmmlu answer generation high school biology, task595 mocha answer generation, task751 svamp subtraction question answering, task1656 gooaq answer generation, task1431 head qa answer generation, task1296 wiki hop question answering, task490 mwsc options generation, task867 mawps multiop question answering, task865 mawps addsub question answering, task1133 xcsr nl commonsense mc classification, task1422 mathqa physics, task1135 xcsr en commonsense mc classification, task054 multirc write correct answer, task1661 super glue classification, task708 mmmlu answer generation high school physics, task1726 mathqa correct answer generation, task664 mmmlu answer generation abstract algebra, task1412 web questions question answering, task002 quoref answer generation, task752 svamp multiplication question answering, task1297 qasc question answering, task692 mmmlu answer generation computer security, task1136 xcsr fr commonsense mc classification, task727 mmmlu answer generation prehistory, task725 mmmlu answer generation nutrition, task104 semeval 2019 task10 closed vocabulary mathematical answer generation, task694 mmmlu answer generation econometrics, task820 protoqa answer generation, task700 mmmlu answer generation high school chemistry, task390 torque text span selection, task1421 mathqa other, task918 coqa answer generation, task309 race answer generation, task247 dream answer generation, task695 mmmlu answer generation electrical engineering, task230 iirc passage classification, task712 mmmlu answer generation high school world history, task731 mmmlu answer generation professional psychology, task596 mocha question generation, task698 mmmlu answer generation global facts, task718 mmmlu answer generation machine learning, task395 persianqa answer generation, task597 cuad answer generation, task339 record answer generation, task835 mathdataset answer generation, task238 iirc answer from passage answer generation, task228 arc answer generation easy, task380 boolq yes no question, task152 tomqa find location easy noise, task754 svamp common-division question answering, task713 mmmlu answer generation human aging, task665 mmmlu answer generation anatomy, task706 mmmlu answer generation high school mathematics, task697 mmmlu answer generation formal logic, task753 svamp addition question answering, task1727 wiqa what is the effect, task1139 xcsr ru commonsense mc classification, task1134 xcsr hi commonsense mc classification, task344 hybridqa answer generation, task165 mcscript question answering commonsense, task1145 xcsr jap commonsense mc classification, task1295 adversarial qa question answering, task239 tweetqa answer generation, task1382 quarel write correct answer, task866 mawps multidiv question answering, task151 tomqa find location easy clean, task591 sciq answer generation, task717 mmmlu answer generation logical fallacies, task726 mmmlu answer generation philosophy, task1419 mathqa gain, task1423 mathqa geometry, task231 iirc link classification, task049 multirc questions needed to answer, task734 mmmlu answer generation sociology, task580 socialiqa answer generation, task736 mmmlu answer generation virology, task729 …
@@
Translation 394 task808 pawsx chinese korean translation, task254 spl translation fi en, task1111 ted translation he it, task988 pib translation oriya english, task650 opus100 ar en translation, task763 emea es lt translation, task1648 opus books en-sv translation, task1263 ted translation pl fa, task1020 pib translation telugu oriya, task913 bianet translation, task1060 pib translation urdu malayalam, task1676 xquad-ca translation, task1098 ted translation ja fa, task984 pib translation marathi gujarati, task1086 pib translation marathi english, task789 pawsx french english translation, task1110 ted translation he gl, task1689 qed amara translation, task787 pawsx korean chinese translation, task1071 pib translation malayalam marathi, task548 alt translation en ch, task1373 newscomm translation, task1023 pib translation english hindi, task1271 ted translation fa it, task1274 ted translation pt en, task552 alt translation en bu, task1040 pib translation punjabi oriya, task1323 open subtitles hi en translation, task1058 pib translation urdu english, task1105 ted translation ar gl, task1353 hind encorp translation en hi, task1085 pib translation english marathi, task1103 ted translation es fa, task784 pawsx korean french translation, task811 pawsx chinese german translation, task1365 opustedtalks translation, task1278 ted translation pt he, task1115 alt ja id translation, task538 alt translation bu en, task786 pawsx korean german translation, task805 pawsx german chinese translation, task1692 qed amara translation, task1690 qed amara translation, task655 bible en fa translation, task1256 ted translation pl en, task977 pib translation oriya urdu, task841 para pdt de en translation, task996 pib translation english bengali, task531 europarl es en translation, task452 opus paracrawl en ig translation, task1250 ted translation it ar, task644 refresd translation, task1248 ted translation it ja, task1034 pib translation hindi gujarati, task1225 ted translation ja he, task997 pib translation bengali oriya, task1127 alt ja th translation, task783 pawsx korean english translation, task1031 pib translation bengali telugu, task560 alt translation en entk, task1000 pib translation tamil malayalam, task252 spl translation en tr, task1650 opus books en-fi translation, task654 bible fa en translation, task802 pawsx german korean translation, task1025 pib translation bengali punjabi, task262 spl translation ja en, task785 pawsx korean spanish translation, task530 europarl en es translation, task1232 ted translation ar es, task799 pawsx spanish chinese translation, task1119 alt fil ja translation, task260 spl translation zh en, task1686 menyo20k translation, task448 opus paracrawl en tl translation, task994 pib translation tamil hindi, task1065 pib translation punjabi telugu, task557 alt translation en ba, task1072 pib translation marathi malayalam, task535 alt translation ch en, task762 emea fr sk translation, task1024 pib translation hindi english, task914 bianet translation, task779 pawsx english spanish translation, task547 alt translation entk en, task1128 alt th ja translation, task537 alt translation th en, task1277 ted translation pt ar, task1124 alt ja lo translation, task1514 flores translation entone, task435 alt en ja translation, task425 hindienglish corpora en hi translation, task1371 newscomm translation, task818 pawsx japanese chinese translation, task873 opus xhosanavy translation xhosa eng, task1240 ted translation gl es, task553 alt translation en ma, task1351 opus100 translation gu en, task999 pib translation malayalam tamil, task438 eng guj parallel corpus en gu translation, task541 alt translation kh en, task1329 open subtitles en hi translation, task1102 ted translation es pl, task661 mizan en fa translation, task1259 ted translation pl ar, task424 hindienglish corpora hi en translation, task793 pawsx french chinese translation, task1005 pib translation malayalam punjabi, task1262 ted translation pl it, task1367 opustedtalks translation, task117 spl translation en de, task1237 ted translation he ar, task1122 alt khm ja translation, task1230 ted translation ar en, task790 pawsx french korean translation, task433 alt hi en translation, task253 spl translation en zh, task1037 pib translation telugu urdu, task840 para pdt en es translation, task982 pib translation tamil bengali, task1009 pib translation bengali hindi, task1062 pib translation marathi bengali, task1218 ted translation en ja, task1113 ted translation he fa, task1691 qed amara translation, task1276 ted translation pt es, task1108 ted translation ar fa, task1070 pib translation urdu bengali, task1244 ted translation gl pl, task1239 ted translation gl ja, task1055 pib translation marathi oriya, task794 pawsx french japanese translation, task1004 pib translation malayalam bengali, task1049 pib translation malayalam telugu, task989 pib translation marathi urdu, task450 opus paracrawl so en translation, task815 pawsx japanese french translation, task1066 pib translation telugu punjabi, task777 pawsx english korean translation, task542 alt translation ja en, task830 poleval2019 mt translation, task1655 mkb translation, task313 europarl en sv translation, task1044 pib translation punjabi gujarati, task1038 pib translation urdu telugu, task1057 pib translation english urdu, task1047 pib translation english telugu, task1258 ted translation pl es, task1001 pib translation gujarati urdu, task1063 pib translation gujarati tamil, task1649 opus books en-no translation, task1282 ted translation pt fa, task983 pib translation gujarati marathi, task261 spl translation es en, task439 eng guj parallel corpus gu en translation, task795 pawsx spanish english translation, task1046 pib translation telugu hindi, task1233 ted translation ar he, task1112 ted translation he pl, task663 global voices en fa translation, task662 global voices fa en translation, task1376 newscomm translation, task258 spl translation fa en, task1029 pib translation marathi punjabi, task986 pib translation oriya hindi, task1067 pib translation bengali gujarati, task604 flores translation entosn, task1224 ted translation ja ar, task250 spl translation en ar, task1242 ted translation gl he, task559 alt translation en fi, task1015 pib translation punjabi tamil, task259 spl translation tr en, task1269 ted translation fa he, task807 pawsx chinese english translation, task809 pawsx chinese french translation, task995 pib translation bengali english, task1093 ted translation en fa, task174 spl translation en ja, task1036 pib translation urdu tamil, task1245 ted translation gl fa, task1081 pib translation hindi marathi, task1249 ted translation it es, task842 para pdt cs en translation, task1006 pib translation punjabi malayalam, task1091 ted translation en it, task1022 pib translation malayalam english, task656 quran en fa translation, task1435 ro sts parallel language translation ro to en, task1219 ted translation en es, task782 pawsx english japanese translation, task1054 pib translation urdu hindi, task1035 pib translation tamil urdu, task554 alt translation en la, task993 pib translation hindi tamil, task540 alt translation la en, task556 alt translation en ja, task1324 open subtitles te en…
Sentence Compression 1 task1340 msr text compression compression
@@
Summarization 16 task1357 xlsum summary generation, task672 amazon and yelp summarization dataset summarization, task1579 gigaword incorrect summarization, task1658 billsum summarization, task618 amazonreview summary text generation, task522 news editorial summary, task1355 sent comp summarization, task589 amazonfood summary text generation, task1553 cnn dailymail summarization, task1572 samsum summary, task1291 multi news summarization, task668 extreme abstract summarization, task1309 amazonreview summary classification, task1499 dstc3 summarization, task1290 xsum summarization, task511 reddit tifu long text summarization
@@
Text to Code 12 task210 logic2text structured text generation, task107 splash question to sql, task077 splash explanation to sql, task076 splash correcting sql mistake, task130 scan structured text generation command action long, task869 cfq mcd1 sql to explanation, task212 logic2text classification, task126 scan structured text generation command action all, task211 logic2text classification, task128 scan structured text generation command action short, task868 cfq mcd1 explanation to sql, task956 leetcode 420 strong password check
@@
Title Generation 19 task1540 parsed pdfs summarization, task1561 clickbait new bg summarization, task769 qed summarization, task1342 amazon us reviews title, task1356 xlsum title generation, task569 recipe nlg text generation, task1161 coda19 title generation, task220 rocstories title classification, task219 rocstories title answer generation, task1586 scifact title generation, task602 wikitext-103 answer generation, task1358 xlsum title generation, task1659 title generation, task418 persent title generation, task743 eurlex summarization, task288 gigaword summarization, task500 scruples anecdotes title generation, task619 ohsumed abstract title generation, task510 reddit tifu title summarization

### v2

@@ Appendix D AI Assistants Using
Question Answering 207 Task 837: viquiquad answer generation, Task 701: mmmlu answer generation high school computer science, Task 1399: obqa answer generation, Task 075: squad1.1 answer generation, Task 724: mmmlu answer generation moral scenarios, Task 666: mmmlu answer generation astronomy, Task 742: lhoestq answer generation frequency, Task 1438: doqa cooking answer generation, Task 863: asdiv multiop question answering, Task 864: asdiv singleop question answering, Task 058: multirc question answering, Task 669: ambigqa answer generation, Task 704: mmmlu answer generation high school government and politics, Task 728: mmmlu answer generation professional accounting, Task 740: lhoestq answer generation quantity, Task 1293: kilt tasks hotpotqa question answering, Task 849: pubmedqa answer generation, Task 1424: mathqa probability, Task 1625: disfl qa asnwer generation, Task 858: inquisitive span detection, Task 723: mmmlu answer generation moral disputes, Task 083: babi t1 single supporting fact answer generation, Task 118: semeval 2019 Task 10: open vocabulary mathematical answer generation, Task 582: naturalquestion answer generation, Task 237: iirc answer from subtext answer generation, Task 714: mmmlu answer generation human sexuality, Task 444: com qa question paraphrases answer generation, Task 720: mmmlu answer generation marketing, Task 332: tellmewhy answer generation, Task 119: semeval 2019 Task 10: geometric mathematical answer generation, Task 310: race classification, Task 1132: xcsr ur commonsense mc classification, Task 702: mmmlu answer generation high school european history, Task 710: mmmlu answer generation high school statistics, Task 870: msmarco answer generation, Task 047: miscellaneous answering science questions, Task 711: mmmlu answer generation high school us history, Task 1286: openbookqa question answering, Task 598: cuad answer generation, Task 685: mmmlu answer generation clinical knowledge, Task 084: babi t1 single supporting fact identify relevant fact, Task 1420: mathqa general, Task 1520: qa srl answer generation, Task 868: mawps singleop question answering, Task 768: qed text span selection, Task 061: ropes answer generation, Task 041: qasc answer generation, Task 144: subjqa question answering, Task 1570: cmrc2018 answer generation, Task 1610: xquad es answer generation, Task 164: mcscript question answering text, Task 703: mmmlu answer generation high school geography, Task 705: mmmlu answer generation high school macroeconomics, Task 1131: xcsr es commonsense mc classification, Task 1130: xcsr vi commonsense mc classification, Task 750: aqua multiple choice answering, Task 473: parsinlu mc classification, Task 385: socialiqa incorrect answer generation, Task 691: mmmlu answer generation college physics, Task 719: mmmlu answer generation management, Task 1327: qa zre answer generation from question, Task 715: mmmlu answer generation international law, Task 737: mmmlu answer generation world religions, Task 010: mctaco answer generation event ordering, Task 741: lhoestq answer generation place, Task 028: drop answer generation, Task 730: mmmlu answer generation professional medicine, Task 491: mwsc answer generation, Task 716: mmmlu answer generation jurisprudence, Task 732: mmmlu answer generation public relations, Task 735: mmmlu answer generation us foreign policy, Task 898: freebase qa answer generation, Task 887: quail answer generation, Task 024: cosmosqa answer generation, Task 1140: xcsr pl commonsense mc classification, Task 225: english language answer generation, Task 1608: xquad en answer generation, Task 170: hotpotqa answer generation, Task 667: mmmlu answer generation business ethics, Task 699: mmmlu answer generation high school biology, Task 595: mocha answer generation, Task 751: svamp subtraction question answering, Task 1656: gooaq answer generation, Task 1431: head qa answer generation, Task 1296: wiki hop question answering, Task 490: mwsc options generation, Task 867: mawps multiop question answering, Task 865: mawps addsub question answering, Task 1133: xcsr nl commonsense mc classification, Task 1422: mathqa physics, Task 1135: xcsr en commonsense mc classification, Task 054: multirc write correct answer, Task 1661: super glue classification, Task 708: mmmlu answer generation high school physics, Task 1726: mathqa correct answer generation, Task 664: mmmlu answer generation abstract algebra, Task 1412: web questions question answering, Task 002: quoref answer generation, Task 752: svamp multiplication question answering, Task 1297: qasc question answering, Task 692: mmmlu answer generation computer security, Task 1136: xcsr fr commonsense mc classification, Task 727: mmmlu answer generation prehistory, Task 725: mmmlu answer generation nutrition, Task 104: semeval 2019 Task 10: closed vocabulary mathematical answer generation, Task 694: mmmlu answer generation econometrics, Task 820: protoqa answer generation, Task 700: mmmlu answer generation high school chemistry, Task 390: torque text span selection, Task 1421: mathqa other, Task 918: coqa answer generation, Task 309: race answer generation, Task 247: dream answer generation, Task 695: mmmlu answer generation electrical engineering, Task 230: iirc passage classification, Task 712: mmmlu answer generation high school world history, Task 731: mmmlu answer generation professional psychology, Task 596: mocha question generation, Task 698: mmmlu answer generation global facts, Task 718: mmmlu answer generation machine learning, Task 395: persianqa answer generation, Task 597: cuad answer generation, Task 339: record answer generation, Task 835: mathdataset answer generation, Task 238: iirc answer from passage answer generation, Task 228: arc answer generation easy, Task 380: boolq yes no question, Task 152: tomqa find location easy noise, Task 754: svamp common-division question answering, Task 713: mmmlu answer generation human aging, Task 665: mmmlu answer generation anatomy, Task 706: mmmlu answer generation high school mathematics, Task 697: mmmlu answer generation formal logic, Task 753: svamp addition question answering, Task 1727: wiqa what is the effect, Task 1139: xcsr ru commonsense mc classification, Task 1134: xcsr hi commonsense mc classification, Task 344: hybridqa answer generation, Task 165: mcscript question answering commonsense, Task 1145: xcsr jap commonsense mc classification, Task 1295: adversarial qa question answering, Task 239: tweetqa answer generation, Task 1382: quarel write correct answer…
@@
Translation 394 Task 808: pawsx chinese korean translation, Task 254: spl translation fi en, Task $1$: ted translation he it, Task 988: pib translation oriya english, Task 650: opus100 ar en translation, Task 763: emea es lt translation, Task 1648: opus books en-sv translation, Task 1263: ted translation pl fa, Task 1020: pib translation telugu oriya, Task 913: bianet translation, Task 1060: pib translation urdu malayalam, Task 1676: xquad-ca translation, Task 1098: ted translation ja fa, Task 984: pib translation marathi gujarati, Task 1086: pib translation marathi english, Task 789: pawsx french english translation, Task 1110: ted translation he gl, Task 1689: qed amara translation, Task 787: pawsx korean chinese translation, Task 1071: pib translation malayalam marathi, Task 548: alt translation en ch, Task 1373: newscomm translation, Task 1023: pib translation english hindi, Task 1271: ted translation fa it, Task 1274: ted translation pt en, Task 552: alt translation en bu, Task 1040: pib translation punjabi oriya, Task 1323: open subtitles hi en translation, Task 1058: pib translation urdu english, Task 1105: ted translation ar gl, Task 1353: hind encorp translation en hi, Task 1085: pib translation english marathi, Task 1103: ted translation es fa, Task 784: pawsx korean french translation, Task 811: pawsx chinese german translation, Task 1365: opustedtalks translation, Task 1278: ted translation pt he, Task 1115: alt ja id translation, Task 538: alt translation bu en, Task 786: pawsx korean german translation, Task 805: pawsx german chinese translation, Task 1692: qed amara translation, Task 1690: qed amara translation, Task 655: bible en fa translation, Task 1256: ted translation pl en, Task 977: pib translation oriya urdu, Task 841: para pdt de en translation, Task 996: pib translation english bengali, Task 531: europarl es en translation, Task 452: opus paracrawl en ig translation, Task 1250: ted translation it ar, Task 644: refresd translation, Task 1248: ted translation it ja, Task 1034: pib translation hindi gujarati, Task 1225: ted translation ja he, Task 997: pib translation bengali oriya, Task 1127: alt ja th translation, Task 783: pawsx korean english translation, Task 1031: pib translation bengali telugu, Task 560: alt translation en entk, Task 1000: pib translation tamil malayalam, Task 252: spl translation en tr, Task 1650: opus books en-fi translation, Task 654: bible fa en translation, Task 802: pawsx german korean translation, Task 1025: pib translation bengali punjabi, Task 262: spl translation ja en, Task 785: pawsx korean spanish translation, Task 530: europarl en es translation, Task 1232: ted translation ar es, Task 799: pawsx spanish chinese translation, Task 1119: alt fil ja translation, Task 260: spl translation zh en, Task 1686: menyo20k translation, Task 448: opus paracrawl en tl translation, Task 994: pib translation tamil hindi, Task 1065: pib translation punjabi telugu, Task 557: alt translation en ba, Task 1072: pib translation marathi malayalam, Task 535: alt translation ch en, Task 762: emea fr sk translation, Task 1024: pib translation hindi english, Task 914: bianet translation, Task 779: pawsx english spanish translation, Task 547: alt translation entk en, Task 1128: alt th ja translation, Task 537: alt translation th en, Task 1277: ted translation pt ar, Task 1124: alt ja lo translation, Task 1514: flores translation entone, Task 435: alt en ja translation, Task 425: hindienglish corpora en hi translation, Task 1371: newscomm translation, Task 818: pawsx japanese chinese translation, Task 873: opus xhosanavy translation xhosa eng, Task 1240: ted translation gl es, Task 553: alt translation en ma, Task 1351: opus100 translation gu en, Task 999: pib translation malayalam tamil, Task 438: eng guj parallel corpus en gu translation, Task 541: alt translation kh en, Task 1329: open subtitles en hi translation, Task 1102: ted translation es pl, Task 661: mizan en fa translation, Task 1259: ted translation pl ar, Task 424: hindienglish corpora hi en translation, Task 793: pawsx french chinese translation, Task 1005: pib translation malayalam punjabi, Task 1262: ted translation pl it, Task 1367: opustedtalks translation, Task 117: spl translation en de, Task 1237: ted translation he ar, Task 1122: alt khm ja translation, Task 1230: ted translation ar en, Task 790: pawsx french korean translation, Task 433: alt hi en translation, Task 253: spl translation en zh, Task 1037: pib translation telugu urdu, Task 840: para pdt en es translation, Task 982: pib translation tamil bengali, Task 1009: pib translation bengali hindi, Task 1062: pib translation marathi bengali, Task 1218: ted translation en ja, Task 1113: ted translation he fa, Task 1691: qed amara translation, Task 1276: ted translation pt es, Task 1108: ted translation ar fa, Task 1070: pib translation urdu bengali, Task 1244: ted translation gl pl, Task 1239: ted translation gl ja, Task 1055: pib translation marathi oriya, Task 794: pawsx french japanese translation, Task 1004: pib translation malayalam bengali, Task 1049: pib translation malayalam telugu, Task 989: pib translation marathi urdu, Task 450: opus paracrawl so en translation, Task 815: pawsx japanese french translation, Task 1066: pib translation telugu punjabi, Task 777: pawsx english korean translation, Task 542: alt translation ja en, Task 830: poleval2019 mt translation, Task 1655: mkb translation, Task 313: europarl en sv translation, Task 1044: pib translation punjabi gujarati, Task 1038: pib translation urdu telugu, Task 1057: pib translation english urdu, Task 1047: pib translation english telugu, Task 1258: ted translation pl es, Task 1001: pib translation gujarati urdu, Task 1063: pib translation gujarati tamil, Task 1649: opus books en-no translation, Task 1282: ted translation pt fa, Task 983: pib translation gujarati marathi, Task 261: spl translation es en, Task 439: eng guj parallel corpus gu en translation, Task 795: pawsx spanish english translation, Task 1046: pib translation telugu hindi, Task 1233: ted translation ar he, Task 1112: ted translation he pl, Task 663: global voices en fa translation, Task 662: global voices fa en translation, Task 1376: newscomm translation, Task 258: spl translation fa en, Task 1029: pib translation marathi punjabi, Task 986: pib translation oriya hindi, Task 1067: pib translation bengali gujarati, Task 604: flores translation entosn, Task 1224: ted translation ja ar, Task 250: spl translation en ar, Task 1242: ted translation gl he, Task 559: alt translation en fi, Task 1015: pib translation punjabi tamil, Task 259: spl translation tr en, Task 1269: ted translation fa he, Task 807: pawsx chinese english translation, Task 809: pawsx chinese french translation, Task 995: pib translation bengali english, Task 1093: ted translation en fa, Task 174: spl translation en ja, Task 1036: pib translation urdu tamil, Task 1245: ted translation gl fa, Task 1081: pib translation hindi marathi, Task 1249: ted translation it es, Task 842: para pdt cs en translation, Task 1006: pib translation punjabi malayalam, Task 1091: ted translation en it, Task 1022: pib translation malayalam english, Task 656: quran en fa translation…
Sentence Compression 1 Task 1340: msr text compression compression
@@
Summarization 16 Task 1357: xlsum summary generation, Task 672: amazon and yelp summarization dataset summarization, Task 1579: gigaword incorrect summarization, Task 1658: billsum summarization, Task 618: amazonreview summary text generation, Task 522: news editorial summary, Task 1355: sent comp summarization, Task 589: amazonfood summary text generation, Task 1553: cnn dailymail summarization, Task 1572: samsum summary, Task 1291: multi news summarization, Task 668: extreme abstract summarization, Task 1309: amazonreview summary classification, Task 1499: dstc3 summarization, Task 1290: xsum summarization, Task 511: reddit tifu long text summarization
@@
Text to Code 12 Task 210: logic2text structured text generation, Task 107: splash question to sql, Task 077: splash explanation to sql, Task 076: splash correcting sql mistake, Task 130: scan structured text generation command action long, Task 869: cfq mcd1 sql to explanation, Task 212: logic2text classification, Task 126: scan structured text generation command action all, Task 211: logic2text classification, Task 128: scan structured text generation command action short, Task 868: cfq mcd1 explanation to sql, Task 956: leetcode 420 strong password check
@@
Title Generation 19 Task 1540: parsed pdfs summarization, Task 1561: clickbait new bg summarization, Task 769: qed summarization, Task 1342: amazon us reviews title, Task 1356: xlsum title generation, Task 569: recipe nlg text generation, Task 1161: coda19 title generation, Task 220: rocstories title classification, Task 219: rocstories title answer generation, Task 1586: scifact title generation, Task 602: wikitext-103 answer generation, Task 1358: xlsum title generation, Task 1659: title generation, Task 418: persent title generation, Task 743: eurlex summarization, Task 288: gigaword summarization, Task 500: scruples anecdotes title generation, Task 619: ohsumed abstract title generation, Task 510: reddit tifu title summarization